# Preserving Multilingual Quality While Tuning Query Encoder on English Only

### v3

@@ 2.2 Datasets
We use MSMARCO Nguyen et al. (2018) Triplets [^4] https://huggingface.co/datasets/sentence-transformers/embedding-training-data/blob/main/msmarco-triplets.jsonl.gz for tuning and evaluation. For evaluating the qualities not targeted by tuning, we use the ARXIV dataset with negatives [^5] https://huggingface.co/datasets/primer-ai/arxiv-negatives, which we made from arxiv (version 173) [^6] https://www.kaggle.com/datasets/Cornell-University/arxiv, and the test subset of the XNLI multilingual dataset [^7] https://huggingface.co/datasets/facebook/xnli Conneau et al. (2018). We also use HOTPOTQA [^8] https://hotpotqa.github.io/ Yang et al. (2018) and SQUAD [^9] https://huggingface.co/datasets/rajpurkar/squad_v2 Rajpurkar et al. (2018, 2016) for confirming some observations (Appendices CD).
@@ 2.3 Tuning and evaluations
Unless otherwise specified, we freeze the text encoder and proceed to fine-tune only the query encoder (fully or partially unfrozen) by contrastive learning on MSMARCO (or narrow ARXIV subsets, Appendix K) with a learning rate of 5e-8, batch size of 14 and the triple margin loss with margin $0.1$. Other details are in Appendix E. In our experiments we considered different settings of freezing, batch size, learning rate, the margin of triplet loss, the stopping criterion, weight decay, scheduling versions and optimizers.
@@ 3.2 Learning rate and adiabatic tuning
Our tentative explanation of adiabatic tuning is as follows: At low learning rates of tuning, the system (the encoder weights) remains in the ’minimum’ region found at pretraining. This ’minimum’ region is probably a wide well with uneven ground; the pretraining happened to terminate at some point inside the well. During tuning, the pretraining weight-space of twin encoder becomes just another surface in a family of surfaces, because of the added dimensions (the difference between the weights of the two encoders). We assume that due to continuity, the ’minimum’ region, even if being reshaped, remains a well as the query encoder weights drift away from the weights of the text encoder. Within this well, improvements of all qualities related to the former, pretraining loss, may be still correlated. But if, at high learning rate, the model is strongly modified at some iteration (i.e. by backpropagation on a particular batch), then it may move away from the well. In Appendix M.3 we present a simple attempt to identify and extend the range of adiabatic tuning.
@@ Appendix A Usage of MSMARCO Triplets
The MSMARCO dataset consists of 499184 samples, with each sample being a tuple given as (query, positives, negatives). The “positives” are the correct answers to the query, and the “negatives” are semantically similar, but incorrect answers. For most samples there is only one positive, but many negatives. For our tuning we simply select the very first positive and the very first negative. Thus, each sample gives one triplet (anchor, positive, negative) for contrastive learning, where the query is taken as an anchor.
@@ B.1 Dataset arxiv-negatives
We make use of triplets created from arxiv because this provides our evaluation with a very different kind of text (compared to MSMARCO), and thus allows us to judge robustness of the improvement. For convenience and reproducibility of creating triplets of different levels of difficulty, we made a dataset arxiv-negatives [^10] https://huggingface.co/datasets/primer-ai/arxiv-negatives. The dataset consists of 253140 samples, each sample is a tuple of two elements:
@@
    (1) An ARXIV paper metadata, including its Id, title and abstract and categories.
@@
    (2) List of 21 Ids of other ARXIV papers. The first 20 Ids are the papers that are ’closest’ to the above paper, and sorted from the most to the least similar; the last 21st Id is an Id of a randomly selected paper (not coinciding with Id of the above paper).
@@
For example, to create triplets of difficulty 14, for each paper given by the first tuple element, we pick up a paper corresponding to 14th Id given in the second tuple element. From the first paper we can create query and positive, and from the second paper - negative. Through this work we used two flavours:
@@
    ’Title’: Query is the title of the first paper, and positive is its abstract; negative is the abstract of the second paper.
@@
    ’First’: Query is the first sentence of the abstract of the first paper; positive is the rest of the abstract; negative is the abstract of the second paper, with its first sentence deleted.
@@ B.2 How is it created?
The above dataset is created from the mirror of arxiv [^11] https://huggingface.co/datasets/arxiv-community/arxiv_dataset [^12] https://www.kaggle.com/datasets/Cornell-University/arxiv (version 173) arxiv-metadata-oai-snapshot.jsonl through the following steps:
@@ Appendix C SQUAD
For using SQUAD dataset, we identified (for each query) the given paragraph sentences containing an answer to the query as positives, and the rest of the sentences as negatives. We left samples having at least 1 positive and 1 negative. On average there is 1.3 positives and 4.2 negatives per a query. For evaluation shown in Table 2 we combined train, validation and test subsets. The results are given also for a version ’SQUAD min 5’, in which we have filtered out queries that had less than 5 candidate sentences.
@@ Appendix D HotpotQA
For using HotpotQA, we combined its train and dev subsets. For each query (’question’) both train and dev subsets contain on average 9.95 passages, of which 2 are always positives. For evaluation shown in Table 2 we filtered out queries that had less than 10 passages, and split the dataset into ’easy’, ’medium’ and ’hard’ subsets accordingly to the HotpotQA labels of the difficulty of the samples.
@@ Appendix H Gains on ARXIV for Different Levels of Difficulty
Through the paper we used the easiest version of triplets in the dataset arxiv-negatives, the version that uses randomly selected negatives. Here in Figure 5 we show, for comparison, the fraction of the errors that untuned original $E5$ embeddings do on the other levels of difficulty, and also the corresponding improvements (by Equation 1) after tuning the query encoder on the MSMARCO with frozen embedding block and our default settings (Section 2.3). The statistical significance of the improvements in Figure 5 is estimated as explained in Appendix I.
The difficulty of intentionally close negatives is much harder, but Figure 5 still shows that performance on ARXIV was mostly improved. We used the easiest triplets version for our evaluations through the paper because it more distinctly indicated the trends in the improvements.
@@ Appendix I Significance Test
The changes in cross-lingual qualities corresponding to the third row (emb, frozen embedding block) of Table 4 are shown in comparison with $E5$ and $E5e$ embeddings in Figures 6 and 7. Note that $E5e$ is not a multilingual embedding model. Having a worse start as a multilingual embedding model, $E5e$ also gets much weaker improvements of its multilingual qualities; it is consistent with our understanding of adiabatic tunings (Section 3.2).
@@
$E5$ tuned on these narrow datasets using our ‘default’ regime (Section 2.3) with frozen embedding block mostly improves the PND (positive-negatives discrepancy fraction) as shown in Table 5. The improvements of these narrow-tuned encoders on individual language pairs, assessed by cosine, are shown in Figures 8 and 9.
@@
Dependency of $L12$ tuning on the batch size is shown in Table 10 (number of samples per epoch is 14000) and in Table 11 (number of batches per epoch is 1000). Observations are somewhat similar to $E5$ (Appendix M.1.2), except that generally $L12$ performs not as good as $E5$ and the batch size 7 turns out bad for $L12$.
@@
In attempting to answer these questions, we considered the largest changes in the layers at different learning rates. One suspect layer, by simple crude measures, is output.dense.weight. Here in Tables 14 and 15 we show why we selected this layer. Our motivation here is based on a simple and crude criteria; more detailed research and understanding may reveal better ways to extend the adiabatic tuning regime.
The gains from the tuning by freezing the layer output.dense.weight (in each transformer block) are shown in Figure 11. In comparison to the default tuning (Figure 2) we can see that the adiabatic regime indeed extends from a learning rate of about 6e-8 (as was in Figure 2) to about 1.3e-7. Thus, freezing of output.dense.weight did help to somewhat extend the adiabatic tuning regime. However, further increase of the learning rate results in worse deterioration for the version with frozen output.dense.weight layer, as can be seen for XNLI starting from the rate 1.4e-7.

### v4

@@ 2.2 Datasets
We use MSMARCO Nguyen et al. (2018) Triplets [^4] https://huggingface.co/datasets/sentence-transformers/embedding-training-data/blob/main/msmarco-triplets.jsonl.gz for tuning and evaluation. For evaluating the qualities not targeted by tuning, we use the ARXIV dataset with negatives [^5] https://huggingface.co/datasets/primer-ai/arxiv-negatives, which we made from arxiv (version 173) [^6] https://huggingface.co/datasets/arxiv-community/arxiv_dataset [^7] https://www.kaggle.com/datasets/Cornell-University/arxiv, and the test subset of the XNLI multilingual dataset [^8] https://huggingface.co/datasets/facebook/xnli Conneau et al. (2018). We also use HOTPOTQA [^9] https://hotpotqa.github.io/ Yang et al. (2018) and SQUAD [^10] https://huggingface.co/datasets/rajpurkar/squad_v2 Rajpurkar et al. (2018, 2016) for confirming some observations (Appendices CD).
@@ 2.3 Tuning and evaluations
Unless otherwise specified, we freeze the text encoder and proceed to fine-tune only the query encoder (fully or partially unfrozen) by contrastive learning on MSMARCO (or on narrow ARXIV subsets, Appendix K) with a learning rate of 5e-8, batch size of 14 and the triple margin loss with margin $0.1$. Other details are in Appendix E. In our experiments we considered different settings of freezing, batch size, learning rate, the margin of triplet loss, the stopping criterion, weight decay, scheduling versions and optimizers.
@@ 3.2 Learning rate and adiabatic tuning
Our tentative explanation of adiabatic tuning is as follows: At low learning rates of tuning, the system (the encoder weights) remains in the ’minimum’ region found at pretraining. This ’minimum’ region is probably a wide well with uneven ground; the pretraining happened to terminate at some point inside the well. During tuning, the pretraining weight-space of twin encoder becomes just another surface in a family of surfaces, because of the added dimensions (the difference between the weights of the two encoders). We assume that due to continuity, the ’minimum’ region, even if being reshaped, remains a well as the query encoder weights drift away from the weights of the text encoder. Within this well, improvements of all qualities related to the former, pretraining loss, may be still correlated. But if, at high learning rate, the model is strongly modified at some iteration (i.e. by backpropagation on a particular batch), then it may move away from the well.
3.3 Extending adiabatic tuning range
In attempting to answer these questions, we have considered the largest changes in the layers at different learning rates. One suspect layer, by simple crude measures, is output.dense.weight. In Appendix M.3 in Tables 14 and 15 we show the most changing layers and the blocks to which they belong. Our motivation here is based on a simple and crude criteria; more detailed research and understanding may reveal better ways to extend the adiabatic tuning regime.
The gains from the tuning by freezing the layer output.dense.weight (in each transformer block) are shown in Figure 3. In comparison to the default tuning (Figure 2) we can see that the adiabatic regime indeed extends from a learning rate of about 6e-8 (as was in Figure 2) to about 1.3e-7. Thus, freezing of output.dense.weight did help to somewhat extend the adiabatic tuning regime. However, this did not improve the gains, and further increase of the learning rate results in worse deterioration for the version with frozen output.dense.weight layer, as can be seen for XNLI starting from the rate 1.4e-7.
Another way of trying to stay longer in the original ’minimum’ region during tuning could be by reducing the inertia of the optimizer. We present a simple attempt in Appendix M.8, but the results are mixed.
@@ Appendix A Usage of MSMARCO Triplets
The MSMARCO dataset consists of 499184 samples, with each sample being a tuple given as (query, positives, negatives). The “positives” are the correct answers to the query, and the “negatives” are semantically similar, but incorrect answers. For most samples, there is only one positive, but many negatives. For our tuning we simply select the very first positive and the very first negative. Thus, each sample gives one triplet (anchor, positive, negative) for contrastive learning, where the query is taken as an anchor.
@@ B.1 Dataset arxiv-negatives
We make use of triplets created from arxiv because this provides our evaluation with a very different kind of text (compared to MSMARCO), and thus allows us to judge the robustness of the improvement. For convenience and reproducibility of creating triplets of different levels of difficulty, we made a dataset arxiv-negatives [^11] https://huggingface.co/datasets/primer-ai/arxiv-negatives.
The dataset consists of 253140 samples, each sample is a tuple of two elements:
@@
    An ARXIV paper metadata, including its Id, title and abstract and categories.
@@
    List of 21 Ids of other ARXIV papers. The first 20 Ids are the papers that are ’closest’ to the above paper, and sorted from the most to the least similar; the last 21st Id is an Id of a randomly selected paper (not coinciding with Id of the above paper).
@@
For example, to create triplets of difficulty 14, for each paper given by the first tuple element, we pick up a paper corresponding to 14th Id given in the second tuple element. From the first paper we can create query and positive, and from the second paper, negative. Through this work we used two flavors:
@@
    ‘Title’: The title of the first paper acts as the query and its abstract as the positive; the negative is then the abstract of the second paper.
@@
    ‘First’: The query is the first sentence of the abstract of the first paper; the positive is the rest of the abstract; the negative is the abstract of the second paper, with its first sentence deleted.
@@ B.2 How is it created?
The above dataset is created from the mirror of arxiv (version 173) arxiv-metadata-oai-snapshot.jsonl through the following steps:
@@ Appendix C SQUAD
For using the SQUAD dataset, we identified (for each query) the given paragraph sentences containing an answer to the query as positives, and the rest of the sentences as negatives. We left samples having at least 1 positive and 1 negative. On average there is 1.3 positives and 4.2 negatives per a query. For the evaluation shown in Table 2 we combined train, validation and test subsets. The results are given also for a version called ‘SQUAD min 5’, in which we have filtered out queries that had less than 5 candidate sentences.
@@ Appendix D HotpotQA
For using HotpotQA, we combined its train and dev subsets. For each query (‘question’) both train and dev subsets contain on average 9.95 passages, of which 2 are always positives. For the evaluation shown in Table 2 we filtered out queries that had less than 10 passages, and split the dataset into ‘easy’, ‘medium’ and ‘hard’ subsets accordingly to the HotpotQA labels of the difficulty of the samples.
@@ Appendix H Gains on ARXIV for Different Levels of Difficulty
Throughout the paper we used the easiest version of triplets in the arxiv-negatives dataset, the version that uses randomly selected negatives. Here in Figure 6 we show, for comparison, the fraction of the errors which occur in the original untuned $E5$ embeddings using the other levels of difficulty, and also the corresponding improvements (by Equation 1) after tuning the query encoder on the MSMARCO with frozen embedding block and our default settings (Section 2.3). The statistical significance of the improvements in Figure 6 is estimated as explained in Appendix I.
The difficulty of intentionally close negatives is much harder, but Figure 6 still shows that performance on ARXIV was mostly improved. We used the easiest triplets version for our evaluations throughout the paper because it more distinctly indicated the trends in the improvements.
@@ Appendix I Significance Test
The changes in cross-lingual qualities corresponding to the third row (emb, frozen embedding block) of Table 4 are shown in comparison with $E5$ and $E5e$ embeddings in Figures 7 and 8. Note that $E5e$ is not a multilingual embedding model. Having a worse start as a multilingual embedding model, $E5e$ also gets much weaker improvements of its multilingual qualities; it is consistent with our understanding of adiabatic tunings (Section 3.2).
@@
$E5$ tuned on these narrow datasets using our ‘default’ regime (Section 2.3) with frozen embedding block mostly improves the PND (positive-negatives discrepancy fraction) as shown in Table 5. The improvements of these narrow-tuned encoders on individual language pairs, assessed by cosine, are shown in Figures 9 and 10.
@@
The dependency of tuning $L12$ using different batch size is shown in Table 10 (number of samples per epoch is 14000) and in Table 11 (number of batches per epoch is 1000). Observations are somewhat similar to $E5$ (Appendix M.1.2), except that generally $L12$ does not perform as well as $E5$ and a batch size of 7 turns out to be bad for $L12$.
@@
In Section 3.3 we showed how the adiabatic tuning range gets extended when the layer output.dense.weight is frozen (in all blocks). The reason for suspecting that this layer is the most responsible for breaking out of the original ‘minimum’ region, is that its maximal weight becomes the highest among all the layers as the learning rate gets closer to the end of the adiabatic range: see Table 14. The maximal relative change of the weights is also achieved by the layer output.dense.weight: see Table 15.
It is a crude adjustment, and freezing this layer in all blocks is probably overkill, but this did help us in extending the adiabatic range (Section 3.3).